7. Inductive reactance on the complex plane is? a. Positive b. Negative e -t otation is? ( 2.112B N +15 8. 250 - 7440 in polaCan you please help me with the questions on this page, preferably the bottom two. Thank you
Answer #1

1) inductive reactance on the complex plane is Positive (option-a) 250 -3440 = 506.063) – 60.39 55061-68 option 3) 25) 40 = 25) ?T= Ri+jWl; = 6ootj 10 (250x2 TX 003) - 600+ j471.238 = 762.93138.145 IT Vslv.m. s) (Ym.s) = ZT 25x52 762.93|38-1459 50004